OpenStudy (seratul):

Need Help With SAT Prep A 36 inch cord is cut into two pieces. of the length of the shorter piece is 1/3 the longer piece, how many inches long is the longer piece?

OpenStudy (wolf1728):

x + (1/3 x) = 36 4/3 x = 36 That's a good start for you.

OpenStudy (seratul):

I got 27

OpenStudy (wolf1728):

That is correct. and the other piece (which you didn't have to calculate) is (1/3) 27 or 9. Also 27 + 9 = 36, the length of the complete cord.